# Understanding Oversmoothing in GNNs as Consensus in Opinion Dynamics

### v1

@@ Appendix A Proof of Propositions
A BIMP model with time-independent bias parameter $\mathbf{B}$, converges to an equilibrium.
@@
$\dot{\mathbf{X}}=-d\mathbf{X}+\operatorname{tanh}\biggl{[}u(\mathbf{A^{\mathrm{a}}}\mathbf{X}+\alpha\mathbf{X})\biggr{]}+\mathbf{X}(0).$ (39)
@@
$\hat{\mathbf{A}}^{\mathrm{a}}=\mathbf{P}\mathbf{A^{\mathrm{a}}}\mathbf{P}^{\top}=\begin{bmatrix}\mathbf{A}_{11}&0&0&0\\ 0&\mathbf{A}_{22}&0&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\dots&\mathbf{A}_{nn}\end{bmatrix},\;\;\hat{\mathbf{X}}=\mathbf{P}\mathbf{X}=\begin{bmatrix}\mathbf{X}_{1}\\ \mathbf{X}_{2}\\ \vdots\\ \mathbf{X}_{n}\end{bmatrix},\;\;\hat{\mathbf{X}}(0)=\mathbf{P}\mathbf{X}(0)=\begin{bmatrix}\mathbf{X}_{1}(0)\\ \mathbf{X}_{2}(0)\\ \vdots\\ \mathbf{X}_{n}(0)\end{bmatrix},$ (40)
@@
$\mathbf{A^{\mathrm{a}}}=\mathbf{P}^{\top}\hat{\mathbf{A}}^{\mathrm{a}}\mathbf{P},\quad\mathbf{X}=\mathbf{P}^{\top}\hat{\mathbf{X}},\quad\mathbf{X}(0)=\mathbf{P}^{\top}\hat{\mathbf{X}}(0).$ (41)
Substituting $\mathbf{A^{\mathrm{a}}}$ and $\mathbf{X}$ with $\hat{A}^{\mathrm{a}}$ and $\hat{\mathbf{X}}$ respectively in Equation (39) yields
$\mathbf{P}^{\top}\dot{\hat{\mathbf{X}}}$ $=-d\mathbf{P}^{\top}\hat{\mathbf{X}}+\operatorname{tanh}\biggl{[}u(\mathbf{P}^{\top}\hat{\mathbf{A}}^{\mathrm{a}}\mathbf{PP^{\top}}\hat{\mathbf{X}}+\alpha\mathbf{P}^{\top}\hat{\mathbf{X}})\biggr{]}+\mathbf{P}^{\top}\hat{\mathbf{X}}(0),$ (42)
$=\mathbf{P}^{\top}(-d\hat{\mathbf{X}})+\mathbf{P}^{\top}\operatorname{tanh}\biggl{[}u(\hat{\mathbf{A}}^{\mathrm{a}}\hat{\mathbf{X}}+\alpha\hat{\mathbf{X}})\biggr{]}+\mathbf{P}^{\top}\hat{\mathbf{X}}(0),$ (43)
@@
$\dot{\hat{\mathbf{X}}}=-d\hat{\mathbf{X}}+\operatorname{tanh}\biggl{[}u(\hat{\mathbf{A}}^{\mathrm{a}}\hat{\mathbf{X}}+\alpha\hat{\mathbf{X}})\biggr{]}+\hat{\mathbf{X}}(0).$ (44)
@@
$\dot{\mathbf{X}}_{m}=-d\mathbf{X}_{m}+\operatorname{tanh}\biggl{[}u(\mathbf{A}_{mm}\mathbf{X}_{m}+\alpha\mathbf{X}_{m})\biggr{]}+\mathbf{X}_{m}(0).$ (45)
@@
  • Cooperative. Since $\operatorname{sech}\in(0,1]$, $\mathbf{A}_{mm}$ are positive matrices (as it is the output of a softmax function) and $\alpha\geq 0$, $\mathbf{J}_{m}(\mathbf{X}_{m})$ is a Metzler matrix in which all the off-diagonal elements are non-negative. This implies that $\mathbf{J}_{m}(\mathbf{X}_{m})$ is cooperative and the subsystem $\dot{\mathbf{X}}_{m}=f_{m}(\mathbf{X}_{m})$ is monotone (Hirsch, 1985).
@@
If the Jacobian for a continuous dynamical system $\dot{x}=f(x,t)$ is cooperative and irreducible, then it approaches the equilibrium for almost every point $x$ whose forward orbit has compact closure (Hirsch, 1985). Since the Jacobian $\mathbf{J}_{m}(\mathbf{X}_{m})$ for each subsystem $\dot{\mathbf{X}}_{m}=f_{m}(\mathbf{X}_{m})$ satisfies this condition, almost every state $\mathbf{X}_{m}$ approaches the equilibrium set. Therefore the dynamical system defined in Equation (39) converges to an equilibrium set. As all trajectories tend towards the equilibrium solution, analyzing the equilibrium behavior is sufficient to understand the underlying dynamics of our BIMP model.
@@
A.3 Proposition 5.7: Reduced order representation of BIMP
When the bias parameter is equal to $\mathbf{B=0}$, the dynamics of BIMP can be approximated by the dynamics of the reduced one-dimensional dynamical equation.
$\dot{y}=-d\,y+\operatorname{tanh}\left[u(\tilde{\alpha}+4)y\right],$ (47)
where $y=\langle\mathbf{x},\mathbf{w}_{\max}\rangle\in\mathbb{R}$, and $\mathbf{w}_{\max}$ is the left dominant eigenvector of $\tilde{\mathbf{A}}$.
Proof.
Leveraging the Lyapunov-Schimit reduction, the BIMP dynamics can be projected onto a one-dimensional critical subspace (Leonard et al., 2024). The BIMP dynamics in Equation (20) can be vectorized following Proposition 5.3 as
$\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\left[u\left(\tilde{\alpha}\mathbf{x}+\tilde{\mathbf{A}}\mathbf{x}\right)\right]+\mathbf{b},$ (48)
$\mathbf{T}\dot{\mathbf{y}}=-d\mathbf{T}\dot{\mathbf{y}}+\operatorname{tanh}\left[u(\tilde{\alpha}\mathbf{T}\mathbf{y}+\tilde{\mathbf{A}}\mathbf{T}\mathbf{y})\right].$ (49)
Multiplying $\mathbf{T}^{-1}$ on both sides gives
$\dot{\mathbf{y}}=-d\dot{\mathbf{y}}+\mathbf{T}^{-1}\operatorname{tanh}\left[u(\tilde{\alpha}\mathbf{T}\mathbf{y}+\tilde{\mathbf{A}}\mathbf{T}\mathbf{y})\right].$ (50)
$\dot{\mathbf{y}}=-d\dot{\mathbf{y}}+\operatorname{tanh}\left[u(\tilde{\alpha}\mathbf{y}+\mathbf{T}^{-1}\tilde{\mathbf{A}}\mathbf{T}\mathbf{y})\right].$ (51)
$\dot{\mathbf{y}}=-d\dot{\mathbf{y}}+\operatorname{tanh}\left[u(\tilde{\alpha}\mathbf{y}+\boldsymbol{\Lambda}\mathbf{y})\right].$ (52)
$\dot{y_{1}}=-d\dot{y_{1}}+\operatorname{tanh}\left[u(\tilde{\alpha}y_{1}+\lambda_{1}y_{1})\right].$ (53)
$\dot{y}=-d\dot{y}+\operatorname{tanh}\left[u(\tilde{\alpha}y+4y)\right],$ (54)
A.4 Proposition 5.9: Formation of dissensus in BIMP
BIMP will not exhibit oversmoothing when the bias parameter $\mathbf{B}$ is time-independent with unique entries.
@@
where $\tilde{\mathbf{a}}_{i}$ is the $i$-th row of $\tilde{\mathbf{A}}$ and $b_{i}$ is the $i$-th element of $\mathbf{b}$. Assume $\mathbf{x}$ converges to consensus such that $x_{1}=x_{2}=\ldots=x_{N_{a}}=\bar{x}$. For any pair $x_{m}$ and $x_{n},m\neq n$ with corresponding bias $b_{m}\neq b_{n}$, their dynamical equations are
@@
due to $\tilde{\mathbf{a}}_{m}$ and $\tilde{\mathbf{a}}_{n}$ being right-stochastic. However, since $b_{m}\neq b_{n}$, the right hand side of Equation (57) and Equation (58) cannot be 00 at the same time. Therefore, by contradiction, consensus cannot be the equilibrium for BIMP if $b_{m}\neq b_{n}$. If $\mathbf{b}$ has unique elements, the equilibrium of the system forms dissensus and avoids oversmoothing.
If there are more than one option in the system (i.e, $\mathbf{A}^{o}\neq\mathbf{0}$), formation of dissensus is still possible since $\mathbf{A}^{\mathrm{o}}$ is also right-stochastic. ∎
A.5 Proposition 5.10 & 5.11: BIMP gradients are well behaved
@@
With increasing depth (i.e, $M\rightarrow\infty$), this repeated multiplication leads to gradient exploding (vanishing) when the components $\nicefrac{{\partial\mathbf{X}^{t}}}{{\partial\mathbf{X}^{t-1}}}>\mathbf{I}$ ($\nicefrac{{\partial\mathbf{X}^{t}}}{{\partial\mathbf{X}^{t-1}}}<\mathbf{I}$). The BIMP model provides an upper and lower bound on gradients in Proposition 5.10 and 5.11 to guarantee exploding or vanish gradients cannot occur.
A.5.1 Proposition 5.10: BIMP gradients don’t explode
Given step-size $\Delta t\ll 1$ and $d\Delta t<1$, BIMP ensures an upper bound for the gradient, preventing it from exploding.
@@
$\mathbf{X}^{t}$ $=\mathbf{X}^{t-1}+\Delta t\biggl{(}-d\mathbf{X}^{t-1}+\tanh\biggl{[}u\biggl{(}\tilde{\alpha}\mathbf{X}^{t-1}+\bigl{(}\mathbf{A}^{\mathrm{a}}+\mathbf{I}\bigr{)}\mathbf{X}^{t-1}\bigl{(}\mathbf{A}^{\mathrm{o\top}}+\mathbf{I}\bigr{)}\biggr{)}\biggr{]}+\mathbf{X}^{0}\biggr{)},$ (67)
@@
$\left\|\frac{\partial\mathbf{x}^{t}}{\partial\mathbf{x}^{t-1}}\right\|_{\infty}$ $\leq\left\|(1-d\Delta t)\mathbf{I}+u\Delta t(\tilde{\alpha}\mathbf{I}+\tilde{\mathbf{A}})\right\|_{\infty},$ (76)
@@
Since $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I})$ from Definition 5.2 and given that $\mathbf{A}^{a}$ and $\mathbf{A}^{o}$ are right-stochastic, it follows that
@@
Consider the third term $\left\|\frac{\partial\mathbf{x}^{1}}{\partial\mathbf{x}^{0}}\right\|_{\infty}$. Since the bias term $\mathbf{x}^{0}$ contributes to the differential defined in Equation (68), it follows that the upper bound can be derived as
@@
which indicates that the gradients are upper bounded regardless of network depth and avoids exploding gradients. ∎
A.5.2 Proposition 5.11: BIMP gradients don’t vanish exponentially
Given step-size $\Delta t\ll 1$ and $d\Delta t<1$, although it still could be small, gradients in BIMP won’t vanish exponentially.
@@
$\mathbf{E}_{t-1}=-d\mathbf{I}+\mathbf{1}\left[\operatorname{sech}^{2}\left(u(\tilde{\alpha}\mathbf{X}^{t-1}+(\mathbf{A}^{a}+\mathbf{I})\mathbf{X}^{t-1}({\mathbf{A}^{o}}^{\top}+\mathbf{I})\right)\right]^{\top}\circ u\biggl{[}\tilde{\alpha}\mathbf{I}+\biggl{(}\mathbf{I}\otimes(\mathbf{A}^{o}+\mathbf{I})\biggr{)}\biggl{(}(\mathbf{A}^{a}+\mathbf{I})\otimes\mathbf{I}\biggr{)}\biggr{]}.$ (102)
@@
$\mathbf{E}_{0}=(1-d)\mathbf{I}+\mathbf{1}\text{vec}\biggl{(}\operatorname{sech}^{2}\left(u(\tilde{\alpha}\mathbf{X}^{t-1}+(\mathbf{A}^{a}+\mathbf{I})\mathbf{X}^{t-1}({\mathbf{A}^{o}}^{\top}+\mathbf{I})\right)\biggl{)}^{\top}\circ u\biggl{[}\tilde{\alpha}\mathbf{I}+\biggl{(}\mathbf{I}\otimes(\mathbf{A}^{o}+\mathbf{I})\biggr{)}\biggl{(}(\mathbf{A}^{a}+\mathbf{I})\otimes\mathbf{I}\biggr{)}\biggr{]}.$ (104)
@@
Appendix B Datasets

### v2

@@ A.2 Behavior-Inspired Message Passing Neural Network
A BIMP model with time-independent input parameter $\mathbf{B}$, converges to an equilibrium.
@@
$\dot{\mathbf{X}}=-d\mathbf{X}+\operatorname{tanh}\biggl{[}u(\mathbf{A^{\mathrm{a}}}\mathbf{X}+\alpha\mathbf{X})\biggr{]}+\mathbf{B}.$ (39)
@@
$\hat{\mathbf{A}}^{\mathrm{a}}$ $=\mathbf{P}\mathbf{A^{\mathrm{a}}}\mathbf{P}^{\top}=\begin{bmatrix}\mathbf{A}_{11}&0&0&0\\ 0&\mathbf{A}_{22}&0&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\dots&\mathbf{A}_{nn}\end{bmatrix},\;\;\;\hat{\mathbf{X}}=\mathbf{P}\mathbf{X}=\begin{bmatrix}\mathbf{X}_{1}\\ \mathbf{X}_{2}\\ \vdots\\ \mathbf{X}_{n}\end{bmatrix},\;\;\;\hat{\mathbf{B}}=\mathbf{P}\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}\\ \mathbf{B}_{2}\\ \vdots\\ \mathbf{B}_{n}\end{bmatrix},$ (40)
@@
$\mathbf{A^{\mathrm{a}}}=\mathbf{P}^{\top}\hat{\mathbf{A}}^{\mathrm{a}}\mathbf{P},\quad\mathbf{X}=\mathbf{P}^{\top}\hat{\mathbf{X}},\quad\mathbf{B}=\mathbf{P}^{\top}\hat{\mathbf{B}}.$ (41)
Substituting $\mathbf{A^{\mathrm{a}}}$ and $\mathbf{X}$ with $\hat{\mathbf{A}}^{\mathrm{a}}$ and $\hat{\mathbf{X}}$ respectively in Equation (39) yields
$\mathbf{P}^{\top}\dot{\hat{\mathbf{X}}}$ $=-d\mathbf{P}^{\top}\hat{\mathbf{X}}+\operatorname{tanh}\biggl{[}u(\mathbf{P}^{\top}\hat{\mathbf{A}}^{\mathrm{a}}\mathbf{PP^{\top}}\hat{\mathbf{X}}+\alpha\mathbf{P}^{\top}\hat{\mathbf{X}})\biggr{]}+\mathbf{P}^{\top}\hat{\mathbf{B}},$ (42)
$=\mathbf{P}^{\top}(-d\hat{\mathbf{X}})+\mathbf{P}^{\top}\operatorname{tanh}\biggl{[}u(\hat{\mathbf{A}}^{\mathrm{a}}\hat{\mathbf{X}}+\alpha\hat{\mathbf{X}})\biggr{]}+\mathbf{P}^{\top}\hat{\mathbf{B}},$ (43)
@@
$\dot{\hat{\mathbf{X}}}=-d\hat{\mathbf{X}}+\operatorname{tanh}\biggl{[}u(\hat{\mathbf{A}}^{\mathrm{a}}\hat{\mathbf{X}}+\alpha\hat{\mathbf{X}})\biggr{]}+\hat{\mathbf{B}}.$ (44)
@@
$\dot{\mathbf{X}}_{m}=-d\mathbf{X}_{m}+\operatorname{tanh}\biggl{[}u(\mathbf{A}_{mm}\mathbf{X}_{m}+\alpha\mathbf{X}_{m})\biggr{]}+\mathbf{B}_{m}.$ (45)
@@
  • Cooperative. Since $\operatorname{sech}\in(0,1]$, $\mathbf{A}_{mm}$ are positive matrices (as it is the output of a softmax function) and $\alpha\geq 0$, $\mathbf{J}_{m}(\mathbf{X}_{m})$ is a Metzler matrix in which all the off-diagonal elements are non-negative. This implies that $\mathbf{J}_{m}(\mathbf{X}_{m})$ is cooperative and the subsystem $\dot{\mathbf{X}}_{m}=f_{m}(\mathbf{X}_{m})$ is monotone [92].
@@
If the Jacobian for a continuous dynamical system $\dot{x}=f(x,t)$ is cooperative and irreducible, then it approaches the equilibrium for almost every point $x$ whose forward orbit has compact closure [92]. Since the Jacobian $\mathbf{J}_{m}(\mathbf{X}_{m})$ for each subsystem $\dot{\mathbf{X}}_{m}=f_{m}(\mathbf{X}_{m})$ satisfies this condition, almost every state $\mathbf{X}_{m}$ approaches the equilibrium set. Therefore the dynamical system defined in Equation (39) converges to an equilibrium set. As all trajectories tend towards the equilibrium solution, analyzing the equilibrium behavior is sufficient to understand the dynamics of BIMP.
@@
A.2.5 Theorem 5.6: Dissensus in BIMP
BIMP will not exhibit oversmoothing when the input parameter $\mathbf{B}$ is time-independent with unique entries.
@@
where $\tilde{\mathbf{a}}_{i}$ is the $i$-th row of $\tilde{\mathbf{A}}$ and $b_{i}$ is the $i$-th element of $\mathbf{b}$. Assume $\mathbf{x}$ converges to consensus such that $x_{1}=x_{2}=\ldots=x_{N_{a}}=\bar{x}$. For any pair $x_{m}$ and $x_{n},m\neq n$ with corresponding input $b_{m}\neq b_{n}$, their dynamical equations are
@@
due to $\tilde{\mathbf{a}}_{m}$ and $\tilde{\mathbf{a}}_{n}$ being right stochastic. However, since $b_{m}\neq b_{n}$, the right hand side of Equation (49) and Equation (50) cannot be 00 at the same time. Therefore, by contradiction, consensus cannot be the equilibrium for BIMP if $b_{m}\neq b_{n}$. If $\mathbf{b}$ has unique elements, the equilibrium of the system forms dissensus and avoids oversmoothing.
If there are more than one option in the system (i.e, $\mathbf{A}^{o}\neq\mathbf{0}$), formation of dissensus is still possible since $\mathbf{A}^{\mathrm{o}}$ is also right stochastic. ∎
A.2.6 Theorem 5.7: BIMP has well behaved gradients
BIMP gradients are upper bounded and do not vanish exponentially.
Proof.
@@
With increasing depth (i.e, $M\rightarrow\infty$), this repeated multiplication leads to gradient exploding (vanishing) when the components $\nicefrac{{\partial\mathbf{X}^{t}}}{{\partial\mathbf{X}^{t-1}}}>\mathbf{I}$ ($\nicefrac{{\partial\mathbf{X}^{t}}}{{\partial\mathbf{X}^{t-1}}}<\mathbf{I}$). The BIMP model provides an upper and lower bound on gradients in Lemma A.3 and A.4 to guarantee exploding or vanish gradients cannot occur. ∎
Lemma A.3.
BIMP gradients are upper bounded when the step-size $\Delta t\ll 1$ and damping term $d<\nicefrac{{1}}{{\Delta t}}$.
@@
$\mathbf{X}^{t}$ $=\mathbf{X}^{t-1}+\Delta t\biggl{(}-d\mathbf{X}^{t-1}+\tanh\biggl{[}u\biggl{(}\tilde{\alpha}\mathbf{X}^{t-1}+\bigl{(}\mathbf{A}^{\mathrm{a}}+\mathbf{I}\bigr{)}\mathbf{X}^{t-1}\bigl{(}\mathbf{A}^{\mathrm{o\top}}+\mathbf{I}\bigr{)}\biggr{)}\biggr{]}+\mathbf{X}^{0}\biggr{)}$ (59)
@@
$\left\|\frac{\partial\mathbf{x}^{t}}{\partial\mathbf{x}^{t-1}}\right\|_{\infty}$ $\leq\left\|(1-d\Delta t)\mathbf{I}+u\Delta t(\tilde{\alpha}\mathbf{I}+\tilde{\mathbf{A}})\right\|_{\infty}$ (68)
@@
Since $\tilde{\mathbf{A}}=(\mathbf{A}^{\mathrm{o}}+\mathbf{I})\otimes(\mathbf{A}^{\mathrm{a}}+\mathbf{I})$ from Definition 5.1 and given that $\mathbf{A}^{a}$ and $\mathbf{A}^{o}$ are right stochastic, it follows that
@@
Consider the third term $\left\|\frac{\partial\mathbf{x}^{1}}{\partial\mathbf{x}^{0}}\right\|_{\infty}$. Since the input term $\mathbf{x}^{0}$ contributes to the differential defined in Equation (60), it follows that the upper bound can be derived as
@@
which indicates that the gradients are upper bounded and avoids exploding gradients. ∎
Lemma A.4.
@@
$\mathbf{E}_{t-1}=-d\mathbf{I}+\mathbf{1}\left[\operatorname{sech}^{2}\left(u(\tilde{\alpha}\mathbf{X}^{t-1}+(\mathbf{A}^{a}+\mathbf{I})\mathbf{X}^{t-1}({\mathbf{A}^{o}}^{\top}+\mathbf{I})\right)\right]^{\top}\\ \circ u\biggl{[}\tilde{\alpha}\mathbf{I}+\biggl{(}\mathbf{I}\otimes(\mathbf{A}^{o}+\mathbf{I})\biggr{)}\biggl{(}(\mathbf{A}^{a}+\mathbf{I})\otimes\mathbf{I}\biggr{)}\biggr{]}.$ (94)
@@
$\mathbf{E}_{0}=(1-d)\mathbf{I}+\mathbf{1}\text{vec}\biggl{(}\operatorname{sech}^{2}\left(u(\tilde{\alpha}\mathbf{X}^{t-1}+(\mathbf{A}^{a}+\mathbf{I})\mathbf{X}^{t-1}({\mathbf{A}^{o}}^{\top}+\mathbf{I})\right)\biggl{)}^{\top}\\ \circ u\biggl{[}\tilde{\alpha}\mathbf{I}+\biggl{(}\mathbf{I}\otimes(\mathbf{A}^{o}+\mathbf{I})\biggr{)}\biggl{(}(\mathbf{A}^{a}+\mathbf{I})\otimes\mathbf{I}\biggr{)}\biggr{]}.$ (96)
@@
Appendix B Other Proofs
We provide additional Lemmas to provide deeper insight into the theoretical properties of our BIMP model.
Lemma B.1 (Expressive capacity of BIMP).
BIMP can model more diverse node feature representations than approaches whose dynamics are equivalent to linear opinion dynamics.
Proof.
Nonlinear systems can exhibit more complex behavior, such as multiple equilibria and bifurcations, than linear systems. Also, many continuous-depth GNNs [26, 48, 49, 50] lack a feature mixing mechanism. In contrast, BIMP introduces the option graph $\mathcal{G}^{\mathrm{o}}$ to enable feature mixing across dimensions, modeling more complex information exchange.
To highlight the contributions of the intrinsic nonlinearity and the option graph $\mathcal{G}^{\mathrm{o}}$, we compare BIMP with linear opinion dynamics as a representative baseline: linear opinion dynamics is a first-order approximation of BIMP without correlated options $\mathbf{A^{\mathrm{o}}}=\mathbf{0}$.
When attention parameter $u=1$, input parameter $\mathbf{B=0}$, and uncorrelated options $\mathbf{A^{\mathrm{o}}}=\mathbf{0}$, the BIMP model has dynamics of the form
When $c=1$, this equation reduces to,
$\dot{\mathbf{X}}=(\mathbf{A}^{\mathrm{a}}-\mathbf{I})\mathbf{X},$ (104)
which is of the same form as linear opinion dynamics in Equation (6). Since linear opinion dynamics has the same form as the first-order approximation of BIMP, we say BIMP has greater expressive capacity. ∎
Lemma B.2 (Expressive capacity of BIMP can degrade).
The BIMP model reduces to a linear system when the attention parameter $u$ is either very small or very large.
Proof.
The degeneration to a linear model occurs under two settings: (1) when $u$ is very small and the nonlinear term evaluates to 0; (2) or when $u$ is very large such that the hyperbolic tangent saturates, and therefore the nonlinear term evaluates to ±$\pm 1$. ∎
To avoid both degenerate cases, we set the attention parameter $u$ at the bifurcation point. Beyond the reasoning provided in Section 5.2, this lemma offers an additional perspective that placing $u$ at the bifurcation point ensures that BIMP operates within the nonlinear regime of Equation (11), thereby preserving its expressive capacity.
Lemma B.3 (Reduced order representation of BIMP).
When the input parameter $\mathbf{B}$ is equal to zero, the dynamics of BIMP can be approximated by the dynamics of the reduced one-dimensional dynamical equation. $\dot{y}=-d\,y+\operatorname{tanh}\left[u(\alpha+3)y\right],$ (105) where $y=\langle\mathbf{x},\mathbf{w}_{\max}\rangle\in\mathbb{R}$, and $\mathbf{w}_{\max}$ is the left dominant eigenvector of $\tilde{\mathbf{A}}$.
Proof.
Leveraging the Lyapunov-Schimit reduction, the BIMP dynamics can be projected onto a one-dimensional critical subspace [35]. The BIMP dynamics in Equation (12) can be vectorized following Lemma 5.2 as
$\dot{\mathbf{x}}=-d\mathbf{x}+\tanh\left[u\left((\alpha-1)\mathbf{x}+\tilde{\mathbf{A}}\mathbf{x}\right)\right]+\mathbf{b},$ (106)
$\mathbf{T}\dot{\mathbf{y}}=-d\mathbf{T}\dot{\mathbf{y}}+\operatorname{tanh}\left[u\biggl{(}(\alpha-1)\mathbf{T}\mathbf{y}+\tilde{\mathbf{A}}\mathbf{T}\mathbf{y}\biggr{)}\right].$ (107)
Multiplying $\mathbf{T}^{-1}$ on both sides gives
$\dot{\mathbf{y}}=-d\dot{\mathbf{y}}+\mathbf{T}^{-1}\operatorname{tanh}\left[u\biggl{(}(\alpha-1)\mathbf{T}\mathbf{y}+\tilde{\mathbf{A}}\mathbf{T}\mathbf{y}\biggr{)}\right].$ (108)
$\dot{\mathbf{y}}=-d\dot{\mathbf{y}}+\operatorname{tanh}\left[u\biggl{(}(\alpha-1)\mathbf{y}+\mathbf{T}^{-1}\tilde{\mathbf{A}}\mathbf{T}\mathbf{y}\biggr{)}\right].$ (109)
$\dot{\mathbf{y}}=-d\dot{\mathbf{y}}+\operatorname{tanh}\left[u\biggl{(}(\alpha-1)\mathbf{y}+\boldsymbol{\Lambda}\mathbf{y}\biggr{)}\right].$ (110)
$\dot{y_{1}}=-d\dot{y_{1}}+\operatorname{tanh}\left[u((\alpha-1)y_{1}+\lambda_{1}y_{1})\right].$ (111)
$\dot{y}=-d\dot{y}+\operatorname{tanh}\left[u(\alpha+3y)\right],$ (112)
Lemma B.4 (Formation of consensus in BIMP).
BIMP exhibits oversmoothing when the input parameter $\mathbf{B}$ is equal to zero.
Proof.
For $\mathbf{x}$ in the neighborhood of the equilibrium $\mathbf{x}=\mathbf{0}$, the Equation (105) in Lemma B.3 is isomorphic to
$\dot{y}=(u(\alpha+3)-d)y-u(\alpha+3)y^{3}.$ (113)
At the bifurcation point $u=\frac{d}{\alpha+3}$, Equation (113) has unique equilibrium $y=0$.
This corresponds to an equilibrium solution of $\mathbf{x}=\mathbf{0}$ in the original system (Equation (36)) which means that $\mathbf{X}=\mathbf{0}$ and all agents form neutral opinions for all options. Since the opinions of all agents have converged, the system has reached consensus (i.e., exhibits oversmoothing). ∎
This lemma indicates that BIMP requires an appropriately chosen input term $\mathbf{B}$ to avoid converging to consensus, as discussed in Theorem 5.6.
Appendix C Datasets
C.1 Homophilic datasets